# Two statistical problems for multivariate mixture distributions

### v1

@@ 3.3.1. A simple example of distances between mixtures
Figure 1. Simulation per Scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively
@@
Figure 2. Boxplot of $500$ replicates of $\textit{MA}_{k}$ calculated between mixtures $F_{1}$ and $F_{2}$ by varying $\eta_{1},\eta_{2}$ and $\eta_{3}$ in the scenarios.
In Figure  1, for better visualization, a simulation per scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. Figure 2 shows how, when the two mixtures $F_{1}$ and $F_{2}$ diverge in location averages, dispersion or weights of the mixture components respectively, the $\textit{MA}_{k}$ average proposed in Algorithm 3.2 captures this separation.
For this purpose, in each scenario, for the selected values of $\eta_{1}$,$\eta_{2}$ and $\eta_{3}$, $500$ values of each mixture are simulated and $\textit{MA}_{k}$ is calculated. This procedure is replicated $500$ times, the Figure 2 shows the boxplots of these $500$ replicates for each case.
When all three parameters, $\eta_{1}$,$\eta_{2}$ and $\eta_{3}$, are zero, both mixtures coincide and the distance between the data simulated by both should be close. As the mixtures are separated, the $\textit{MA}_{k}$ values show an increasing behaviour.
3.3.2. An example of distances between mixtures for quality clustering
An interesting application of Algorithm  3.2 is to measure the quality of a clustering procedure, as discussed previously. In this example we illustrate how the classification error is captured by the distance between mixtures.
Assume that the cluster number $\ell$ is known a priori. We define $\psi_{i}=j$ if the true label of observation $X_{i}$ is $C_{j}$ with $i=1,2,\ldots,n$ and $j=1,2,\ldots,\ell$. Let $\widehat{\psi}_{i}$ be the labels assigned by the clustering procedure, and denote by $\Pi_{\ell}$ the family of permutations of $\{1,2,\ldots,\ell\}$.
The prediction error is given by the minimum between all permutations of the assigned labels and the true labels, that is,
$\min_{\pi\in\Pi_{\ell}}\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{\left\{\psi_{i}\neq\pi\left(\widehat{\psi}_{i}\right)\right\}}.$
On the other hand, $\textit{MA}_{k}$ is calculated between the true mixture and the one generated by the clustering procedure, where $\widehat{\lambda}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{\left\{\widehat{\psi}_{i}=j\right\}}$.
We consider $\ell=2$ and generate a sample of size $n=500$ from the mixture of two bivariate Gaussians distributions, with parameters
$\lambda$ $:=(0.3,0.7),\ \mu_{1}:=(1,-1)\ \quad\text{and}\quad\mu_{2}=(-2,2),$
$\Sigma_{1}$ $:=\begin{pmatrix}1&0.5\\ 0.5&2\end{pmatrix}\quad\text{and}\quad\Sigma_{2}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}.$
The clustering methods implemented are $k$-means, see [19], $k$-medoids (PAM), [27], Spectral clustering, see [34], Hierarchical Ward, see [35] and Mean Shift, [4]. These algorithms were implemented in the R language with the packages stats, cluster, hclust, kernlab and LPCM.
[…]
Figure 3. Estimation of the deviations of the distances obtained for the two mixtures in Scenario 1 when $\eta=1$, as the number of projections $k$ increases.
[…]
4. A new estimator for Gaussian mixtures
4.1. Introduction
There is a large literature on this subject. See for instance the well-known work in [21], the recent work [7] and the references therein. The article [21] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}–f_{j}|\geq\epsilon$. In [7], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before Theorem 4.1 below.
4.2. Description of the algorithm
Let $P$ be a Gaussian mixture on $\mathbb{R}^{d}$, with density
$\sum_{j=1}^{n}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{d/2}}\exp\Bigl{(}\frac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{n}\in[0,1]$ with $\sum_{j=1}^{n}\lambda_{j}=1$, where $\mu_{1},\dots,\mu_{n}\in\mathbb{R}^{d}$, and where $\Sigma_{1},\dots,\Sigma_{n}$ are $d\times d$ positive-definite matrices.
We write $\Lambda:=(\lambda_{1},\dots,\lambda_{n})$ and $\bm{\mu}:=(\mu_{1},\ldots,\mu_{n})$ and $\bm{\Sigma}:=(\Sigma_{1},\ldots,\Sigma_{n})$. Let $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$ denote the set of parameters.
Given a unit vector $u\in\mathbb{R}^{d}$ and $\theta\in\Theta$, we denote by $P_{\langle u\rangle}(\theta)$ the projection of the Gaussian mixture with parameters $\theta$ in the direction $u$. Given an iid sample of the Gaussian mixture $X_{1},\ldots,X_{N}$, let $P_{N,\langle u\rangle}$ be the projection of the empirical measure in the direction $u$.
Fix $k\geq(1/2)(\ell+m-1)(d^{2}+d-2)+1$, and let $u_{1},\ldots,u_{k}$ be iid random directions in $\mathbb{R}^{d}$.
We perform a two-step sequential estimation algorithm. In the first step, the vector $\Lambda$ is estimated. In the second step, the remaining parameters of the mixture are estimated.
Step 1: As is well known, if the data is projected using a linear transformation and a Gaussian mixture model is fitted on the original data, then the density on the projection subspace is again a mixture of Gaussian variables with covariance matrices and mean vectors obtained by a linear transformation of the parameters estimated in the original scale of the data.
For each random direction $u\in\mathbb{R}^{d}$, we can fit a univariate Gaussian mixture of the projected data as follows. We consider the model given by $\sum_{j=1}^{n}\lambda_{j}N_{u}(\upsilon_{j},\sigma^{2}_{j})$, where $\upsilon:=(\upsilon_{1},\dots,\upsilon_{n})$ and $\sigma^{2}:=(\sigma^{2}_{1},\dots,\sigma_{n}^{2})$ are the means and the variances of the projected distributions, given by $\upsilon_{j}:=\langle u,\mu_{j}\rangle,\ \sigma^{2}_{j}:=\langle u,\Sigma_{j}u\rangle\ (j=1,\dots n).$ Thus the distribution of the projected data is given by $\sum_{j=1}^{n}\lambda_{j}N\Bigl{(}\langle u,\mu_{j}\rangle,\langle u,\Sigma_{j}u\rangle\Bigr{)}$, and its characteristic function is
$\psi_{u}(t)=\psi_{u}(t,\upsilon,\sigma^{2},\Lambda):=\sum_{j=1}^{n}\lambda_{j}\exp[i\langle u,\mu_{j}\rangle t+\textstyle\frac{1}{2}\langle u,\Sigma_{j}u\rangle t^{2}].$
We denote by $\widehat{\psi}_{u}$ the empirical characteristic function (ECF) in the direction $u$, that is, $\widehat{\psi}_{u}(t)=\frac{1}{N}\sum_{j=1}^{N}\exp(i\langle u,X_{j}\rangle t)$. As in [33] and [37], we use the ECF procedure to estimate the parameters of mixtures of normal distributions. Roughly speaking, the objective is to minimize the distance between the characteristic function of the model and its empirical counterpart in the data. In the example, we consider a set of fixed points of the grid $t_{1},\dots,t_{m}$ where both functions are compared, see [33].
Let
$Z_{u}$ $:=\Bigl{(}\operatorname{Re}\psi_{u}(t_{1}),\ldots,\operatorname{Re}\psi_{u}(t_{m}),\operatorname{Im}\psi_{u}(t_{1}),\dots,\operatorname{Im}\psi_{u}(t_{m})\Bigr{)}^{t},$
$F_{u}(\upsilon,\sigma^{2},\Lambda)$ $:=\Bigl{(}\operatorname{Re}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u}(t_{m}),\operatorname{Im}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u}(t_{m})\Bigr{)}^{t},$
where $\operatorname{Re}z$ and $\operatorname{Im}z$ denote the real and imaginary parts of the complex number $z$ respectively. We define the estimator $(\widehat{\upsilon}_{u,(1)},\widehat{\sigma}^{2}_{u,(1)},\widehat{\Lambda}_{u,(1)})$ as the values of the parameters that minimize the product
$\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2})\bigr{)}^{t}W\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2})\bigr{)},$
where $W$ is some weighting matrix defined so as to ensure that the sample orthogonality conditions are made as close as possible to zero. The optimal matrix $W$ can be estimated by means of a heteroskedasticity and consistent autocorrelation (HAC) matrix such as the one proposed in [22]. This is just the generalized moment estimator, see [15], where the moment conditions are of the form $h(t,u,X_{j},\upsilon,\sigma^{2})=\exp(it\langle u,X_{j}\rangle)-\psi_{u}(t,\upsilon,\sigma^{2})$, and obviously $E(h(t_{i},u,X_{j},\upsilon,\sigma^{2},\Lambda))=0$ for all $i=1,\dots,m.$
The procedure is replicated in $k$ random directions $u_{1},\dots,u_{k}$, and the estimator of $\Lambda$ is $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widehat{\Lambda}_{u_{r},(1)}$.
Step 2: Replacing $\Lambda$ by its estimated value $\widehat{\Lambda}_{(1)}$, we repeat Step 1, but only for the parameters $(\upsilon_{u},\sigma^{2}_{u})$. The estimators
$\widehat{\upsilon}_{u_{r},(2)}$ $:=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),n}),\,\,\widehat{\sigma}^{2}_{u_{r},(2)}:=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),n}),$
are obtained for $r=1,\dots,k$. Then the estimators for the mixture parameters can be obtained by least squares:
(3) $\widehat{\mu}_{j}:=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr{)}^{2},$
(4) $\widehat{\Sigma}_{j}:=\operatorname*{argmin}_{\Sigma_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2},$
for all $j=1,\dots,n$.
Remark.
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr{|},\ \ \widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr{|},$
for all $j=1,2,\dots,n$.
(iii) We have to make sure that the estimators $\Sigma_{j}$ are positive semi-definite matrices, by using semidefinite optimization as follows (see for instance [17]).
Let $\mathcal{S}^{d}$ denote the space of real symmetric $d\times d$ matrices, equipped with the usual inner product $\langle S_{1},S_{2}\rangle:=\operatorname{trace}(S_{1}^{t}S_{2})$, and let $\mathcal{S}^{+,d}$ be the closed convex cone consisting of positive semi-definite matrices. Let $\Pi_{\mathcal{S}^{+,d}}$ be the projection of $\mathcal{S}^{d}$ onto $\mathcal{S}^{+,d}$, given by $\Pi_{\mathcal{S}^{+,d}}(S):=\operatorname*{argmin}_{S^{\prime}\in\mathcal{S}^{+,d}}\|S^{\prime}-S\|\quad(S\in\mathcal{S}^{d})$. It turns out that, if the spectral decomposition of $S\in\mathcal{S}^{d}$ is given by $S=U\Gamma U^{t}$, where $\Gamma=\operatorname{diag}(\gamma_{1},\dots,\gamma_{d})$, then $\Pi_{\mathcal{S}^{+,d}}(S)=U\Gamma^{+}U^{t}$, where $\Gamma^{+}=\operatorname{diag}(\gamma_{1}^{+},\dots,\gamma_{d}^{+})$ and $\gamma_{j}^{+}=\max(\gamma_{j},0)\leavevmode\nobreak\ (j=1,\dots,d)$.
If we want a strictly positive-definite matrix, then we can add a term $\delta I_{d}$, where $I_{d}$ is the $d\times d$ identity matrix and $\delta$ is a very small positive constant.
(iv) Lastly, we note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
4.3. A consistency result
The results in [33] and [37] that we will use below are based on a general theorem in [16], where the only assumption is that the characteristic function is differentiable. More precisely, it is assumed that the regular case holds, namely that $I_{n}(t)$ can be differentiated under the integral sign.
Theorem 4.1.
Under the weak assumption that,
[…]
Proof.
First observe that, from the results in [33] and [37] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [33, Proposition 1], where strong consistency and asymptotic normality are established.
for the true values of $\mu_{j}$ and $\Sigma_{j}\leavevmode\nobreak\ (j=1,\dots n)$, we have
[…]
and
[…]
for all $j=1,\dots,n$. This concludes the proof. ∎
4.4. Simulations
4.4.1. Example 1.
$F:=\lambda_{1}\textrm{t}(\eta,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\textrm{t}(\eta,\mu_{2},\Sigma_{2}),$
with
$\mu_{1}=(1,0),\quad\mu_{2}=(5,0),\quad\Sigma_{1}=\begin{pmatrix}1&0\\ 0&1/2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}1/2&0\\ 0&1\end{pmatrix},\quad\lambda_{1}=0.3\textrm{ and }\eta=4.$
Figure 5. Depiction of $1000$ simulations of the mixture $F$.
In this Example, we present a small comparative study of the estimation derived by the random projections method (RP) with respect to the expectation-maximization method for mixtures of t-distributions (we denote EM-st). This last method, developed in [24], is implemented by means of the function fmmst of the EMMIXuskew package of the R language. In this case the skew parameter was prefixed to 00.
Table 1. Averages values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 1 0 1 1/2 0
2 5 0 1/2 1 0
RP 1 0.99 -0.01 0.99 0.42 -0.01
sd (0.10) (0.10) (0.16) (0.15) (0.08)
2 4.99 0.014 0.55 0.67 -0.01
sd (0.13) (0.12) (0.06) (0.21) (0.08)
EM-st 1 1.22 0.07 0.89 0.62 -0.01
sd (0.15) (0.16) (0.14) (0.13) (0.09)
2 4.56 -0.10 0.51 1.15 -0.01
sd (0.10) (0.25) (0.07) (0.15) (0.07)
Figure 6. Comparison between the EM-st and RP methods using the $L^{2}$-distance between the parameters.
4.4.2. Example 2
$F=\lambda_{1}N(\mu_{1},\Sigma_{1})+(1-\lambda_{1}-\gamma)N(\mu_{2},\Sigma_{2})+\gamma\mathcal{C}a\left(\mu_{3}\right),$
with
$\mu_{1}=(1,0),\quad\mu_{2}=(5,0),\quad\mu_{3}=(3,0),\quad\Sigma_{1}=\begin{pmatrix}4&0\\ 0&2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}2&0\\ 0&4\end{pmatrix},\quad\lambda_{1}=0.3\textrm{ and }\gamma=0.05.$
Figure 7. Comparison between the Robust EM and RP Boxplots when outliers are added, taking $\mu_{3}=(3,0)$.
4.5. An example with real data
Figure 8. Hexagonal binning for bivariate data. The plane is tessellated by a regular grid of hexagons (cells). Then the hexagons are plotted using an intensity of color proportional to the number of points falling in each cell. Higher socio-economic level for smaller values on the $x$-axis, better mathematical ability for larger values on the $y$-axis.
$\widehat{\Sigma}_{1}=\begin{pmatrix}0.28&0.05\\ 0.05&0.45\end{pmatrix},\quad\widehat{\Sigma}_{2}=\begin{pmatrix}0.16&0.07\\ 0.07&0.40\end{pmatrix}.$
These values are similar to those obtained in the mixture estimation.
5. Conclusions and future work
We consider two important statistical problems: that of testing and measuring the agreement between two different random partitions for cluster analysis, and that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections.
@@ 5. Conclusions and future work
We propose a new estimator for multivariate mixture distributions based on one-dimensional projections and their respective characteristic functions. We derive the strong consistency of them and in a small simulation study we compare with the more classical EM-type estimators, where we observe a very good behaviour.
The relationship between school performance and students’ socioeconomic and cultural status is analyzed in a real data example. We find two different patterns that are possible, explained by the type of institution that the students attend : public or private.
Rates of convergence and the asymptotic distribution of the estimators remain as open problems that are not addressed in the present manuscript.
The number of directions $k$ into which we project is not a smoothing parameter that should be chosen in an optimal way. In practice we suggest to use a larger value for $k$ (without increasing too much the computational time) than the one provided by Theorem 2.1 which corresponds to the case when the mixtures are known, whereas we are dealing with empirical distributions which are close to the unknown underlying distributions but not exactly equal. Depending on the problem, among the directions chosen at random it may appear some directions which are close to bad ones. Increasing the value of $k$ solves the problem. In our experience after a reasonable value the performance stabilizes. We illustrate this fact in Figure 3 .
Also some more computationally oriented work, which includes different mixtures of distributions, sample sizes and number of clusters, is also not included in the present manuscript.

### v3

@@ 4.3 Simulations
Fig. 5: Simulation per Scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively
@@
Fig. 6: Boxplot of $500$ replicates of $\textit{MA}_{k}$ (Right panels) and $D_{k}$ (Left panels), with $k=100$, calculated between mixtures $F_{1}$ and $F_{2}$ by varying $\eta_{1},\eta_{2}$ and $\eta_{3}$ in the scenarios.
In Figure 5, for improved visualization, one representative simulation per scenario is displayed, corresponding to $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. The maximum aggregation measure $D_{k}$ quantifies the largest observed separation between the mixtures $F_{1}$ and $F_{2}$ across the $k$ nearest-neighbor comparisons. This measure is particularly sensitive to pronounced discrepancies (whether in location, dispersion, or component weights), thereby providing a robust indicator of distinct distributional features.
Figure 6 presents, for each scenario, boxplots of the $500$ replicates of $D_{k}$, obtained by simulating 500 observations from each mixture and repeating the procedure 500 times. When $\eta_{1}=\eta_{2}=\eta_{3}=0$, the mixtures coincide and $D_{k}$ attains minimal values. As the mixtures diverge, $D_{k}$ exhibits a clear increasing trend, faithfully reflecting the growing separation.
The same analysis was carried out using the average aggregation measure $\textit{MA}_{k}$, obtaining similar results, as shown in the right column of Figure 6.
5 Conclusions and future work
The article begins with a review of random-projection methods in statistics.
Then we consider two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions, and that of measuring the agreement between two different random partitions for cluster analysis.
We propose a new estimator for multivariate mixture distributions based on one-dimensional projections and their respective characteristic functions. We derive the strong consistency of them and, in a small simulation study, we compare with the more classical EM-type estimators, where we observe a very good behaviour.
@@ 5 Conclusions and future work
The relationship between school performance and students’ socioeconomic and cultural status is analyzed in a real-data example. We find two different patterns that are possible, explained by the type of institution that the students attend : public or private.
The number of directions $k$ into which we project is not a smoothing parameter that should be chosen in an optimal way. In practice we suggest to use a larger value for $k$ (without increasing too much the computational time) than the one provided by Theorem 5, which corresponds to the case when the mixtures are known, whereas we are dealing with empirical distributions which are close to the unknown underlying distributions but not exactly equal. Depending on the problem, among the directions chosen at random it may appear some directions which are close to bad ones. Increasing the value of $k$ solves the problem. In our experience after a reasonable value the performance stabilizes.
Some more computationally oriented work, which includes different mixtures of distributions, sample sizes and number of clusters, has been omitted from the present manuscript. Also, rates of convergence and the asymptotic distribution of the estimators remain as interesting open problems to be addressed in future work.
Appendix A Algorithm 1: Methodological Aspects
This appendix provides a detailed account of the steps in Algorithm 1.
A.1 Step 1: Estimation of mixture weights
For each random direction $u$, the projected distribution is
$\sum_{j=1}^{n}\lambda_{j}N\!\Bigl{(}\langle u,\mu_{j}\rangle,\langle u,\Sigma_{j}u\rangle\Bigr{)},$
with characteristic function
$\psi_{u}(t)=\sum_{j=1}^{n}\lambda_{j}\exp\!\left(i\langle u,\mu_{j}\rangle t-\tfrac{1}{2}\langle u,\Sigma_{j}u\rangle t^{2}\right).$
Its empirical counterpart is
$\widehat{\psi}_{u}(t)=\frac{1}{N}\sum_{j=1}^{N}\exp(i\langle u,X_{j}\rangle t).$
Let $t_{1},\dots,t_{m}$ be fixed grid points. Define
$Z_{u}$ $:=\bigl{(}\operatorname{Re}\psi_{u}(t_{1}),\dots,\operatorname{Re}\psi_{u}(t_{m}),\operatorname{Im}\psi_{u}(t_{1}),\dots,\operatorname{Im}\psi_{u}(t_{m})\bigr{)}^{T},$
$F_{u}(\upsilon,\sigma^{2},\Lambda)$ $:=\bigl{(}\operatorname{Re}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u}(t_{m}),\operatorname{Im}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u}(t_{m})\bigr{)}^{T}.$
The estimator $(\widehat{\upsilon}_{u,(1)},\widehat{\sigma}^{2}_{u,(1)},\widehat{\Lambda}_{u,(1)})$ minimizes
$\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2},\Lambda)\bigr{)}^{T}W\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2},\Lambda)\bigr{)},$
where $W$ is a weighting matrix, estimated for instance using HAC methods [58]. This corresponds to a generalized method-of-moments estimator [42]. The procedure is replicated in $k$ directions $u_{1},\dots,u_{k}$, and
$\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widehat{\Lambda}_{u_{r},(1)}.$
A.2 Step 2: Estimation of means and covariances
Fixing $\Lambda=\widehat{\Lambda}_{(1)}$, repeat the estimation in each $u_{r}$ for $(\upsilon,\sigma^{2})$. Denote the resulting estimators as $\widehat{\upsilon}_{u_{r},(2)}=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),n})$ and $\widehat{\sigma}^{2}_{u_{r},(2)}=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),n})$.
Then reconstruct the original parameters by solving
$\widehat{\mu}_{j}$ $=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr{)}^{2},$ (6)
$\widehat{\Sigma}_{j}$ $=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2},$ (7)
for $j=1,\dots,n$, with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.